On November 1, 2018, Howell Company purchased 1,000 of the $1,000 face value, 9% bonds of Ramsey, Incorporated, for $1,052,500, which includes accrued interest of $15,000. The bonds, which mature on January 1, 2023, pay interest semiannually on March 1 and September 1. Assuming that Howell uses the straight-line method of amortization and that the bonds are appropriately classified as available-for-sale, the net carrying value of the bonds should be shown on Howell's December 31, 2018, balance sheet at

a. $1,037,500.

b. $1,036,000.

c. $1,052,500.

d. $1,000,000.
Answer #1
calculation of net carrying value of bond
par value of bond (1000*1000) = 1000000
carrying value 1 nov = (1052500-15000) = 1037500
purchase cost of bond including interest = 1052500
premium on bond (1037500-1000000) = 37500
maturity period -remaining = 50
amortization of premium = 37500*2/50 = 1500
net carrying value of bond on balance sheet (1037500-1500)
net carrying value of bond on balance sheet = 1036000
Option "b" is CORRECT